# Semantic Refinement with LLMs for Graph Representations

### v1

@@ Appendix C Structural Features Used for Description Construction
Beyond triangular motifs, the square clustering coefficient characterizes the prevalence of four-node (quadrilateral) structures around a node. It reflects higher-order local dependencies and complementary structural patterns that are not captured by triangle-based clustering alone (zhang2008clustering).
@@ Appendix D Pseudo Code
Algorithm 1 summarizes the overall DAS pipeline. DAS first constructs a structure-aware textual summary for each node and (optionally) concatenates it with the raw node text to initialize descriptions. It then runs a closed-loop refinement for $T$ iterations: (i) encode current descriptions and train the pre-defined GNN, (ii) store semantic–structural–predictive states in a memory buffer, (iii) retrieve an in-graph support set for each node from the buffer, and (iv) refine all node descriptions using an LLM conditioned on the retrieved exemplars. After $T$ iterations, the refined descriptions are used to train the final GNN classifier for evaluation.
Algorithm 1 DAS Pipeline
1:Graph $G=(\mathcal{V},\mathcal{E})$, optional node texts $\{r_{v}\}$, training labels on $\mathcal{V}_{\mathrm{train}}$, fixed-backbone GNN $g_{\theta}$, LLM $\mathcal{M}$, iterations $T$, support size $K$
2:Final descriptions $\{d_{v}^{(T)}\}$ and trained classifier $g_{\theta}^{\ast}$
@@
11:   $d_{v}^{(0)}\leftarrow[\,r_{v}\,\|\,t_{v}^{\text{struct}}\,]$
@@
13:   $d_{v}^{(0)}\leftarrow t_{v}^{\text{struct}}$
@@
16:for $t=0$ to T1T-1 do
17:  Encode $\{d_{v}^{(t)}\}_{v\in\mathcal{V}}$ into features $\{\mathbf{x}_{v}^{(t)}\}$
18:  Train / update $g_{\theta}$ on labeled nodes $\mathcal{V}_{\mathrm{train}}$
19:  Obtain predictions $\{\mathbf{p}_{v}^{(t)}\}_{v\in\mathcal{V}}$
20:  Build memory $\mathcal{B}^{(t)}=\{(d_{v}^{(t)},\mathbf{s}_{v},\mathbf{p}_{v}^{(t)})\}_{v\in\mathcal{V}}$
21:  for all $v\in\mathcal{V}$ do
22:   Retrieve support sets $\mathcal{S}_{v}^{(t)}\leftarrow Retrieve(\mathcal{B}^{(t)},K)$
23:   Refine description $d_{v}^{(t+1)}\leftarrow\mathcal{M}\big(d_{v}^{(t)},\mathcal{S}_{v}^{(t)}\big)$
24:
25:Step 4: Final Training
26:Encode $\{d_{v}^{(T)}\}$ into features and train final GNN $g_{\theta}^{\ast}$
27:
28:return $\{d_{v}^{(T)}\},g_{\theta}^{\ast}$
@@ Appendix E Time Complexity
Let $n=|\mathcal{V}|$ and $m=|\mathcal{E}|$. We summarize the cost of DAS at a high level to clarify that the seemingly complex memory loop adds only modest overhead beyond the GNN and LLM calls. We assume bounded description/prompt lengths, fixed embedding dimensions, and fixed support size $K$.
@@
Per-iteration cost (one refinement round).
Each round consists of three components:
$C_{\text{iter}}(n,m)$ $=\underbrace{C_{\text{gnn}}(n,m)+nC_{\text{enc}}+\mathcal{O}(n)}_{\text{encode + GNN + memory update}}$
@@
$C_{\text{enc}}$ is the description encoding cost, $C_{\text{sent}}$ is the sentence embedding cost used in retrieval, and $C_{\text{llm}}$ is the cost of one LLM call (dominated by prompt+generation tokens). Importantly, the memory construction/update is only $\mathcal{O}(n)$ per round; it does not introduce additional message passing or graph traversal beyond the fixed GNN backbone.
@@
After $T$ refinement rounds, we perform one final encoding and GNN training/evaluation on $\{d_{v}^{(T)}\}$:
$C_{\text{total}}$ $=C_{\text{pre}}(G)+T\,C_{\text{iter}}(n,m)$
$\quad+\big(C_{\text{gnn}}(n,m)+nC_{\text{enc}}\big).$
@@
With bounded text length and fixed $K$, the main costs are (i) the GNN training/inference term $C_{\text{gnn}}(n,m)$, (ii) the retrieval term $\mathcal{O}(n^{2})$ under brute-force similarity computation, and (iii) the LLM term $nC_{\text{llm}}$. All other components introduced by DAS (buffer writes, prompt assembly) are linear in $n$.
@@ Appendix F Further Analysis
F.1 Can DAS Provide Better Initialization for Downstream Learning?
We include an additional pretrain–finetune experiment as a diagnostic study to evaluate a key mechanism of DAS. Specifically, we examine whether node semantics refined by DAS on a source graph provide a better initialization for downstream learning on a related target graph. In this setting, a model is first trained on the source graph and then fine-tuned on the target graph under the high-label split, using a shared textual feature space for all methods. Table 6 reports the results, where Raw Text, TAPE, KEA, and TANS serve as baselines, and DAS is evaluated using the same backbone and sentence encoder. Rather than framing this setting as a standalone transfer learning benchmark, the goal is to probe whether feedback-driven semantic refinement yields node descriptions that are more reusable and better aligned with subsequent optimization on a different but related citation network.
Cora $\to$ Pubmed Pubmed $\to$ Cora
Raw Feat.
   + SVD $70.39\pm 6.12$ $70.48\pm 3.71$
Raw Text $75.77\pm 2.96$ $79.62\pm 2.04$
   + TAPE $75.60\pm 2.39$ $79.25\pm 2.06$
   + KEA $75.25\pm 2.50$ $79.59\pm 1.61$
   + TANS $76.14\pm 2.28$ $80.05\pm 1.74$
    + DAS (Ours) $76.23\pm 2.73$ $80.59\pm 1.32$
Table 6: Results of the pretrain–finetune study probing the quality of refined node semantics as initialization for downstream learning.
F.2 Role of the Semantic–Structural Trade-off
We analyze the role of the trade-off parameter $\alpha$ in the joint similarity score $S(v,u)=\alpha\,\mathrm{sim}_{t}(v,u)+(1-\alpha)\,\mathrm{sim}_{s}(v,u)$, which controls how semantic and structural cues are balanced during exemplar retrieval. This analysis probes whether retrieval based on a single modality is sufficient, or whether effective refinement requires joint alignment.
Figure 4 reports results for Structure-only retrieval ($\alpha=0$), Text-only retrieval ($\alpha=1$), and intermediate values. On the text-attributed Cora graph, text-only retrieval performs competitively, reflecting the strong semantic signal in raw node texts, while structure-only retrieval is weaker. In contrast, on the text-free USA graph, both extremes degrade performance, indicating that neither semantic nor structural similarity alone provides reliable in-context guidance. Sweeping $\alpha\in\{0.1,0.3,0.5,0.7,0.9\}$ further reveals a smooth performance landscape. Higher $\alpha$ values are preferred on Cora, whereas lower values yield better results on USA, suggesting that the optimal balance shifts with the underlying structure–semantics regime. Overall, these results show that joint semantic–structural retrieval is critical for stable refinement, while DAS remains robust to moderate variations in $\alpha$.
Figure 4: Hyper-parameter sensitivity on $\alpha$.
F.3 Computational Cost
DAS relies on a large language model (LLM) to iteratively regenerate node descriptions, and its computational cost is therefore dominated by LLM inference. Specifically, the cost scales with (i) the number of nodes whose descriptions are refined and (ii) the number of refinement iterations. In each refinement round, DAS issues one LLM call per node, resulting in a total of
$\text{LLM calls}=N\times T,$ (28)
where $N$ denotes the number of refined nodes and $T$ the number of refinement iterations.
We report LLM usage in terms of both API calls and token consumption. As an illustrative example, refining N=16N{=}16 nodes for a single iteration (T=1T{=}1) results in $16$ calls to the :contentReference[oaicite:0]index=0 API, consuming 41,69841{,}698 input tokens and 3,4373{,}437 output tokens in total. For larger graphs or additional refinement rounds, the number of calls and token usage increase linearly with $N$ and $T$. In practice, our experiments indicate that a small number of refinement iterations (e.g., $T\leq 3$) is sufficient to capture most of the performance gains, offering a favorable trade-off between accuracy and computational cost.
@@ Appendix G Case Study Results
We present the success cases in Table 7 and failure cases in Table 8.
@@
Dataset: Cora Node: 1103 True label: Theory Pred @ iter 2: Theory Original (default text): Bias, variance and prediction error for classification rules. We…tain a bootstrap estimate of the error of a "bagged" classifier. Refined (iter 2): This paper delves into the relationship between bias, variance, and prediction error in classification frameworks, drawing on Efron’s (1978) seminal work. It meticulously dissects prediction error into its fundamental components, elucidating their effects on classifier performance. The authors introduce a bootstrap-based estimate for the error associated with "bagged" classifiers, demonstrating how bagging and ensemble methods can significantly reduce prediction error. By providing a deeper theoretical understanding and practical tools, this work contributes to improving the accuracy and robustness of classification models.
@@
Dataset: Cora Node: 1148 True label: Theory Pred @ iter 2: Theory Original (default text): Error-based and entropy-based discretization of continuous… with respect to C4.5 and Naive-Bayesian classifiers on datas… Refined (iter 2): This paper presents a detailed comparison of error-based and entropy-based discretization techniques for machine learning classification tasks. It explores how discretization affects model performance and emphasizes the effectiveness of the MDL heuristic in guiding entropy-based discretization. The authors evaluate the methods using classifiers such as the C4.5 decision tree algorithm and Naive Bayes, demonstrating that the entropy-based MDL heuristic generally outperforms error minimization approaches. The study offers both theoretical insights and experimental validation, providing practical guidance for selecting discretization strategies that enhance classification accuracy.
@@
Dataset: Europe-Airports Node: 333 True label: Medium-High Activity Pred @ iter 2: Medium-High Activity Original (default text): Given a node from a airport network graph, where the node …etweenness Centrality" is 0.0000, ranked at 268 among 399 nodes. Refined (iter 2): This airport node is part of a network comprising 399 airports and displays graph characteristics that suggest medium-high activity. It has a Square Clustering Coefficient of 0.3393 (30th rank) and a Clustering Coefficient of 0.6364 (47th rank), indicating strong local clustering and well-connected neighboring airports. However, its degree is 8 (308th rank), showing it has relatively few direct connections compared to more central hubs. The node’s Betweenness Centrality is 0.0000 (268th rank), suggesting it does not serve as a key connector between different parts of the network. Overall, this airport exhibits strong local connectivity but limited global influence, consistent with a medium-high activity airport.
@@
Table 7: Good refinement case studies (text-rich + text-free). For each node, we show the original (default_text) description and the final refined description (iter 2), together with the ground-truth label and the final predicted label. These examples illustrate successful refinement patterns: (i) in text-rich graphs (Cora), the rewrite adds or sharpens topic-discriminative technical cues consistent with the true class (e.g., theory/MDL/entropy-based discretization); (ii) in text-free airport graphs, the rewrite converts raw topological attributes into a coherent role/activity narrative that matches the correct activity bin.
@@
Dataset: Cora Node: 392 True label: Genetic Algorithms Pred @ iter 2: Reinforcement Learning Original (default text): The Canonical Distortion Measure in Feature Space …n. We establish some PAC-like sample complexity bounds …tion problems. … Refined (iter 2): This paper presents the Canonical Distortion Measure (CDM) as an innovative distortion metric designed for improved pattern comparison in machine learning tasks. It addresses shortcomings of traditional distortion measures and provides a theoretical framework with PAC-like bounds on sample complexity. The authors also validate CDM through experimental evaluation, demonstrating its practical relevance. Additionally, the study highlights CDM’s applicability in feature space representation, offering enhanced performance in classification and pattern recognition problems. Overall, the work contributes to both theoretical understanding and practical implementation of distortion measures in machine learning.
Dataset: Cora Node: 714 True label: Genetic Algorithms Pred @ iter 2: Case Based Original (default text): Constructing Fuzzy Graphs from Examples: Methods t… locally independent fuzzy rules … Refined (iter 2): This paper presents a novel and efficient algorithm for constructing fuzzy graphs from datasets by generating locally independent fuzzy rules. The method emphasizes computational efficiency and simplicity, enabling it to model uncertain or complex relationships in the data. It provides a user-friendly framework for representing structured knowledge through fuzzy graph structures, supporting applications in pattern recognition and knowledge representation. The algorithm offers a practical approach to capturing relationships and uncertainties inherent in real-world data.
Dataset: Europe-Airports Node: 322 True label: High Activity Pred @ iter 2: Medium-Low Activity Original (default text): Given a node from a airport network graph, where the node …ss Centrality" is 0.3909, ranked at 55 among 399 nodes. Refined (iter 2): This airport node is part of a network consisting of 399 airports and exhibits high local clustering, as shown by its clustering coefficient of 0.8125 (33rd rank) and square clustering coefficient of 0.4286 (77th rank). Despite this, its degree of 11 (260th rank) indicates relatively few direct connections, limiting its role as a major hub. Its closeness centrality rank of 55th suggests moderate overall reachability in the network. The node’s betweenness centrality is relatively low compared to top connectors, implying it does not frequently lie on shortest paths between airports. Overall, these characteristics suggest that while the airport is well-integrated locally, it has limited global influence and fits a medium-low activity profile.
@@
Appendix H Implementation Details
For each dataset, backbone, and method we perform a random search over architecture and optimization hyperparameters. The candidate values are: hidden dimension {8, 16, 32, 64, 128, 256}, number of layers {1, 2, 3}, normalization layer in {none, batchnorm}, learning rate {5e-2, 1e-2, 5e-3, 1e-3}, weight decay {0.0, 5e-5, 1e-4, 5e-4}, and dropout rate {0.0, 0.1, 0.5, 0.8}. For each configuration we choose the setting that achieves the best validation accuracy. DAS introduces additional hyperparameters: the number of refinement iterations $T$, entropy threshold $\tau$, and the support-set size $K$ used in exemplar retrieval. We fix $K=10$, and $\tau=0.5$ for all datasets. We use $T=3$ refinement iterations on all datasets except Pubmed, where we set $T=1$.
Appendix I Prompts

### v2

@@ Appendix C Structural Features Used for Description Construction
Beyond triangular motifs, the square clustering coefficient characterizes the prevalence of four-node (quadrilateral) structures around a node. It reflects higher-order local dependencies and complementary structural patterns that are not captured by triangle-based clustering alone (Zhang et al., 2008).
@@ Appendix D Pseudo Code
Algorithm 1 summarizes the overall GES pipeline. GES first constructs a structure-aware textual summary for each node and (optionally) concatenates it with the raw node text to initialize descriptions. It then trains a fixed GNN on the initial descriptions to populate a model-conditioned memory, retrieves in-graph exemplar sets for each node, and refines all node descriptions using an LLM conditioned on the retrieved exemplars. The refined descriptions are used to train the final GNN classifier for evaluation.
Algorithm 1 GES Pipeline
1:Graph $G=(\mathcal{V},\mathcal{E})$, optional node texts $\{r_{v}\}$, training labels on $\mathcal{V}_{\mathrm{train}}$, fixed-backbone GNN $g_{\theta}$, LLM $\mathcal{M}$, support size $K$
2:Final descriptions $\{d_{v}^{\prime}\}$ and trained classifier $g_{\theta}^{\ast}$
@@
11:   $d_{v}\leftarrow[\,r_{v}\,\|\,t_{v}^{\text{struct}}\,]$
@@
13:   $d_{v}\leftarrow t_{v}^{\text{struct}}$
@@
16:Encode $\{d_{v}\}_{v\in\mathcal{V}}$ into features $\{\mathbf{x}_{v}\}$
17:Train $g_{\theta}$ on labeled nodes $\mathcal{V}_{\mathrm{train}}$
18:Obtain predictions $\{\mathbf{p}_{v}\}_{v\in\mathcal{V}}$
19:Build memory $\mathcal{B}=\{(d_{v},\mathbf{s}_{v},\mathbf{p}_{v})\}_{v\in\mathcal{V}}$
20:for all $v\in\mathcal{V}$ do
21:  Retrieve support set $\mathcal{S}_{v}\leftarrow\textsc{Retrieve}(\mathcal{B},K)$
22:  Refine description $d_{v}^{\prime}\leftarrow\mathcal{M}(d_{v},\mathcal{S}_{v})$
23:
24:Step 4: Final Training
25:Encode $\{d_{v}^{\prime}\}$ into features and train final GNN $g_{\theta}^{\ast}$
26:
27:return $\{d_{v}^{\prime}\},g_{\theta}^{\ast}$
@@ Appendix E Time Complexity
Let $n=|\mathcal{V}|$ and $m=|\mathcal{E}|$. We summarize the cost of GES at a high level to clarify that the memory-based retrieval adds only modest overhead beyond the GNN and LLM calls. We assume bounded description/prompt lengths, fixed embedding dimensions, and fixed support size $K$.
@@
Refinement cost.
The single-pass refinement consists of three components:
$C_{\text{refine}}(n,m)$ $=\underbrace{C_{\text{gnn}}(n,m)+nC_{\text{enc}}+\mathcal{O}(n)}_{\text{encode + GNN + memory construction}}$
@@
$C_{\text{enc}}$ is the description encoding cost, $C_{\text{sent}}$ is the sentence embedding cost used in retrieval, and $C_{\text{llm}}$ is the cost of one LLM call (dominated by prompt+generation tokens). The memory construction is only $\mathcal{O}(n)$; it does not introduce additional message passing or graph traversal beyond the fixed GNN backbone.
@@
After refinement, we perform one final encoding and GNN training/evaluation on $\{d_{v}^{\prime}\}$:
$C_{\text{total}}=C_{\text{pre}}(G)+C_{\text{refine}}(n,m)+C_{\text{gnn}}(n,m)+nC_{\text{enc}}.$
@@
With bounded text length and fixed $K$, the main costs are (i) the GNN training/inference term $C_{\text{gnn}}(n,m)$, (ii) the retrieval term $\mathcal{O}(n^{2})$ under brute-force similarity computation, and (iii) the LLM term $nC_{\text{llm}}$. All other components introduced by GES (memory writes, prompt assembly) are linear in $n$.
@@ Appendix F Further Analysis
F.1 Embedding Discriminability
Because GES descriptions are conditioned on exemplars from structurally and semantically similar nodes, the LLM rewrites naturally emphasize class-consistent terminology. We quantify this effect by measuring the discriminability gap—the difference between average same-class and different-class cosine similarities in the MiniLM embedding space—and the inter/intra cluster ratio, both computed before GNN training. Figure 5 reports these metrics across all methods. GES achieves the highest discriminability gap on both Cora (+17.0% over TANS) and Pubmed (+35.9%), and the highest inter/intra ratio on the text-free USA graph (+26.4%).
Figure 5: Embedding discriminability before GNN training. GES produces the most class-separable embeddings.
Metric validation.
To confirm that the discriminability gap is a meaningful predictor of downstream performance, we compute correlations across all method–dataset combinations. The Pearson correlation between discriminability gap and test accuracy is $r=0.89$ ($p<0.0001$), and the Spearman rank correlation is $\rho=0.76$ ($p=0.0006$). Per-dataset, the correlation is near-perfect on Pubmed ($r=0.99$, $p=0.012$) and weak on Cora ($r=0.05$, $p=0.94$), consistent with the observation that text-rich graphs are less sensitive to description quality. These results validate the discriminability gap as a reliable proxy for the quality of refined node descriptions.
@@ Appendix G Case Study Results
Why GES Succeeds: Semantic Sharpening and Role Abstraction.
GES is most effective when exemplar-guided refinement sharpens task-discriminative evidence already latent in the input representations. On text-attributed graphs, successful refinement makes class-consistent technical cues more explicit, leading to reduced predictive entropy and corrections toward the true label. On text-free graphs, GES succeeds when raw topological statistics are reorganized into a coherent semantic role interpretation (e.g., distinguishing regionally embedded nodes from global connectors based on clustering and betweenness). In both cases, refinement aligns node semantics more closely with the structure–semantics regime of the graph, enabling the classifier to make more confident and accurate predictions. Representative examples are provided in Table 8.
When GES Fails: Drift and Over-Confidence.
Failure cases reveal inherent limitations of LLM-based semantic refinement. One common failure mode is label drift, where rewrites improve fluency without introducing additional discriminative evidence, causing predictions to shift toward a semantically adjacent but incorrect class. Another failure mode is over-confidence, in which refinement reduces predictive entropy while the prediction itself remains incorrect. On text-free graphs, we additionally observe occasional attribute drift, where numeric structural attributes are subtly altered during rewriting, raising faithfulness concerns even when predictive accuracy improves. These failures highlight that effective refinement depends on maintaining a tight coupling between generated semantics and the underlying structural evidence. See Table 9 for examples.
@@
Dataset: Cora Node: 1103 True label: Theory Pred: Theory Original (default text): Bias, variance and prediction error for classification rules. We…tain a bootstrap estimate of the error of a ”bagged” classifier. Refined: This paper delves into the relationship between bias, variance, and prediction error in classification frameworks, drawing on Efron’s (1978) seminal work. It meticulously dissects prediction error into its fundamental components, elucidating their effects on classifier performance. The authors introduce a bootstrap-based estimate for the error associated with ”bagged” classifiers, demonstrating how bagging and ensemble methods can significantly reduce prediction error. By providing a deeper theoretical understanding and practical tools, this work contributes to improving the accuracy and robustness of classification models.
@@
Dataset: Cora Node: 1148 True label: Theory Pred: Theory Original (default text): Error-based and entropy-based discretization of continuous… with respect to C4.5 and Naive-Bayesian classifiers on datas… Refined: This paper presents a detailed comparison of error-based and entropy-based discretization techniques for machine learning classification tasks. It explores how discretization affects model performance and emphasizes the effectiveness of the MDL heuristic in guiding entropy-based discretization. The authors evaluate the methods using classifiers such as the C4.5 decision tree algorithm and Naive Bayes, demonstrating that the entropy-based MDL heuristic generally outperforms error minimization approaches. The study offers both theoretical insights and experimental validation, providing practical guidance for selecting discretization strategies that enhance classification accuracy.
@@
Dataset: Europe-Airports Node: 333 True label: Medium-High Activity Pred: Medium-High Activity Original (default text): Given a node from a airport network graph, where the node …etweenness Centrality” is 0.0000, ranked at 268 among 399 nodes. Refined: This airport node is part of a network comprising 399 airports and displays graph characteristics that suggest medium-high activity. It has a Square Clustering Coefficient of 0.3393 (30th rank) and a Clustering Coefficient of 0.6364 (47th rank), indicating strong local clustering and well-connected neighboring airports. However, its degree is 8 (308th rank), showing it has relatively few direct connections compared to more central hubs. The node’s Betweenness Centrality is 0.0000 (268th rank), suggesting it does not serve as a key connector between different parts of the network. Overall, this airport exhibits strong local connectivity but limited global influence, consistent with a medium-high activity airport.
@@
Table 8: Good refinement case studies (text-rich + text-free). For each node, we show the original (default_text) description and the refined description, together with the ground-truth label and the predicted label. These examples illustrate successful refinement patterns: (i) in text-rich graphs (Cora), the rewrite adds or sharpens topic-discriminative technical cues consistent with the true class (e.g., theory/MDL/entropy-based discretization); (ii) in text-free airport graphs, the rewrite converts raw topological attributes into a coherent role/activity narrative that matches the correct activity bin.
@@
Dataset: Cora Node: 392 True label: Genetic Algorithms Pred: Reinforcement Learning Original (default text): The Canonical Distortion Measure in Feature Space …n. We establish some PAC-like sample complexity bounds …tion problems. … Refined: This paper presents the Canonical Distortion Measure (CDM) as an innovative distortion metric designed for improved pattern comparison in machine learning tasks. It addresses shortcomings of traditional distortion measures and provides a theoretical framework with PAC-like bounds on sample complexity. The authors also validate CDM through experimental evaluation, demonstrating its practical relevance. Additionally, the study highlights CDM’s applicability in feature space representation, offering enhanced performance in classification and pattern recognition problems. Overall, the work contributes to both theoretical understanding and practical implementation of distortion measures in machine learning.
Dataset: Cora Node: 714 True label: Genetic Algorithms Pred: Case Based Original (default text): Constructing Fuzzy Graphs from Examples: Methods t… locally independent fuzzy rules … Refined: This paper presents a novel and efficient algorithm for constructing fuzzy graphs from datasets by generating locally independent fuzzy rules. The method emphasizes computational efficiency and simplicity, enabling it to model uncertain or complex relationships in the data. It provides a user-friendly framework for representing structured knowledge through fuzzy graph structures, supporting applications in pattern recognition and knowledge representation. The algorithm offers a practical approach to capturing relationships and uncertainties inherent in real-world data.
Dataset: Europe-Airports Node: 322 True label: High Activity Pred Medium-Low Activity Original (default text): Given a node from a airport network graph, where the node …ss Centrality” is 0.3909, ranked at 55 among 399 nodes. Refined: This airport node is part of a network consisting of 399 airports and exhibits high local clustering, as shown by its clustering coefficient of 0.8125 (33rd rank) and square clustering coefficient of 0.4286 (77th rank). Despite this, its degree of 11 (260th rank) indicates relatively few direct connections, limiting its role as a major hub. Its closeness centrality rank of 55th suggests moderate overall reachability in the network. The node’s betweenness centrality is relatively low compared to top connectors, implying it does not frequently lie on shortest paths between airports. Overall, these characteristics suggest that while the airport is well-integrated locally, it has limited global influence and fits a medium-low activity profile.
@@
Appendix H Dataset Statistics
Table 10 summarizes the five benchmark graphs used in our experiments. Cora and Pubmed are text-attributed citation networks where each node carries a title and abstract; USA, Europe, and Brazil are text-free airport graphs where node identity is determined entirely by structural role. The datasets span a wide range of scales (131–19,717 nodes) and structure–semantics regimes, providing a comprehensive testbed for evaluating data-centric semantic refinement.
Nodes Edges Classes Graph Types
Cora 2,708 10,556 7 Text-attribute
Pubmed 19,717 88,648 3 Text-attribute
USA 1,190 28,388 4 Text-free
Europe 399 12,385 4 Text-free
Brazil 131 2,137 4 Text-free
Table 10: Dataset statistics used in our experiments.
Appendix I Implementation Details
For each dataset, backbone, and method we perform a random search over architecture and optimization hyperparameters. The candidate values are: hidden dimension {8, 16, 32, 64, 128, 256}, number of layers {1, 2, 3}, normalization layer in {none, batchnorm}, learning rate {5e-2, 1e-2, 5e-3, 1e-3}, weight decay {0.0, 5e-5, 1e-4, 5e-4}, and dropout rate {0.0, 0.1, 0.5, 0.8}. For each configuration we choose the setting that achieves the best validation accuracy. GES introduces two additional hyperparameters: the entropy threshold $\tau$ and the support-set size $K$ used in exemplar retrieval. For citation graphs, We fix $K=10$ and $\tau=0.5$ for all datasets.
Appendix J Prompts